OpenStudy (anonymous):

How do you find f(x) and g(x) of y=(2/x^2)+3

OpenStudy (anonymous):

And how do you represent it as f(g(x))

OpenStudy (anonymous):

I don't think that would work.

OpenStudy (anonymous):

did you want to decompose y=(2/x^2) + 3 into two functions f and g so that f(g(x)) = (2/x^2) + 3 ???

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

i'm pretty sure what i had earlier was correct... but since i erased it, what do you think it is???

OpenStudy (anonymous):

I checked it, it wasn't correct.

OpenStudy (anonymous):

ARen't there multiple possibilities?

OpenStudy (anonymous):

yes... there are other ways...

OpenStudy (anonymous):

Then why is it asking for a specific answer, is it?

OpenStudy (anonymous):

Sorry, new to this.

OpenStudy (anonymous):

here is another one: f(x) = (2/x^2) + 3 g(x) = x

OpenStudy (anonymous):

That seems extremely easy, won't it work for all of them, except instead of f(x) bing a specific number, it is the equation that is in need of solving?

OpenStudy (anonymous):

being*

OpenStudy (anonymous):

the point is, the decomposition is not unique.... idk why they ask for a specific decomposition... unless they stated some specifics.....

OpenStudy (anonymous):

They just asked for f(x) and g(x). Sorry.

OpenStudy (anonymous):

But thanks!

OpenStudy (anonymous):

:D

OpenStudy (anonymous):

i'll put my original answer down again (in case you missed it): \(\large f(x)=\frac{2}{x}+3 \) \(\large g(x)=x^2 \)

OpenStudy (anonymous):

thanks for the medal... :)

OpenStudy (anonymous):

No prblem

OpenStudy (anonymous):

And once again, thanks a lot!

OpenStudy (anonymous):

yw.. :)